OpenStudy (anonymous):

Solve this system. Please state answer as an ordered pair. 16x-10y=10 -8x-6y=6

OpenStudy (anonymous):

ok ok first step.

OpenStudy (anonymous):

Remember this has the same concept as the last one.

OpenStudy (anonymous):

Solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then for y

OpenStudy (anonymous):

-8x - 6y = 6 -8x = 6 + 6y right so far??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now divide by -8.

OpenStudy (anonymous):

(6 +6y)/8 ? like this?

OpenStudy (anonymous):

The equation would be would be x=-6/8y-6/8, so yes

OpenStudy (anonymous):

Then plug that in for x in the original equation.

OpenStudy (anonymous):

alright so how will it be after that?

OpenStudy (anonymous):

It would be 16(-6/8y-6/8)-10y=10

OpenStudy (anonymous):

Then: -12y-12-10y=10 -2y=22 y=-11 and then plug the y back into the original equation

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

16x-10(-11)=10 16x+110=10 16x=-100 x=25/4

OpenStudy (anonymous):

Coordinates are (25/4,-11)

OpenStudy (anonymous):

Does this help @Keridavis

OpenStudy (anonymous):

Yes thank you.